OpenStudy (anonymous):

Hallie is trying to win the grand prize on a game show. Should she try her luck by spinning a wheel with 6 equal sections labeled from 1 to 6 and hope she gets a 5, or should she roll two number cubes and hope she gets the same number on both cubes?

OpenStudy (anonymous):

p1= 1/6 p2= 1/6, because the first die gives any number and the second has a 1/6 chance of matching it choose either and pray